### v2

@@ Appendix B Analysis of Multi-Patch Prediction
Unlike task-specific, non-autoregressive models such as TFT (Lim et al., 2021) and Informer (Zhou et al., 2021), which predict all future time points simultaneously, their architecture confines them to forecasting a fixed length identical to the training sequence, precluding variable-length predictions. Conversely, TSFMs typically employ an iterative process: they predict a segment of the sequence, append it to the historical data, and then forecast the subsequent segment, thereby enabling autoregressive predictions of arbitrary length. To mitigate the cumulative errors inherent in this autoregressive methodology, Kairos introduces forecast tokens designed to predict multiple patches within each autoregressive step.
In this section, we present an analysis of the multi-patch prediction strategy introduced in Section 3.4. During the training phase, we experimented with a range of forecast patch numbers, specifically $J=\{1,2,4,6,12\}$. The corresponding evaluation results on the GIFT-Eval benchmark are illustrated in Figure 5. Our observations reveal that when $J=1$, which corresponds to the conventional approach of predicting a single patch (Ansari et al., 2024; Liu et al., 2024c; Das et al., 2024), the model achieves optimal performance in short-term forecasting. However, this method necessitates multiple iterations of autoregressive prediction, leading to a significant degradation in performance for medium- and long-term forecasting.
@@
Appendix C Analysis of Inference Speed
@@
To provide a comprehensive view of our model’s efficiency, we benchmarked the single-batch inference speed of Kairos against several state-of-the-art models. The experimental setup involved an input sequence length of 2048 and a prediction horizon of 96 on a single NVIDIA TITAN RTX GPU. For TTM-Advanced, its maximum supported input length of 1536 was used.
@@
As shown in Table 4, the inference time of Kairos is on the same order of magnitude as other highly efficient models like ChronosBolt and Moirai. This demonstrates that Kairos maintains competitive computational efficiency while delivering the superior forecasting accuracy detailed in the main paper. Notably, it is significantly faster than models like Time-MoE, TimesFM, and especially Chronos.
@@
Table 4: Comparison of single-batch inference speeds. All models were tested with an input length of 2048 and an output length of 96, with the exception of TTM-Advanced (*), which used its maximum input length of 1536. Times are averaged per batch and reported in seconds.
Model Inference Time (s)
TTM-Advanced* 0.009
Timer-XL 0.014
ChronosBolt-Base 0.055
Kairos-Base (Ours) 0.061
Moirai-Large 0.070
Sundial-Base 0.099
Time-MoE-Large 0.148
TimesFM-2.0 0.202
Chronos-Large 4.901
Toto-Base 13.717
Appendix D Implementation Details
D.1 Training Details
D.1.1 Model Configurations
@@
D.1.2 Loss Function
To better accommodate varied forecast horizons, and following the methodology of ElasTST (Zhang et al., 2024), we build upon the standard quantile loss by assigning distinct weights to each timestep, such that earlier predictions are given greater importance. The model’s parameters are optimized by minimizing the quantile loss with weight decay, formulated as:
$\mathcal{L}=\frac{1}{B}\sum_{i=1}^{B}\sum_{t=1}^{H}\frac{1}{K}\sum_{k=1}^{K}\omega(t)L_{\alpha_{k}}(y_{i,t},q_{i,t}(\alpha_{k})),$ (8)
$\omega(t)=\frac{1}{H}(\ln(H)-\ln(t)),\quad L_{\alpha}(y,q)=(\alpha-\mathbf{1}_{\{y<q\}})(y-q),$ (9)
where $B,H,K$ are the batch size, forecast horizon, and number of quantiles, respectively; $y_{i,t}$ is the ground truth value, $q_{i,t}(\alpha_{k})$ is the $k$-th quantile forecast, $\omega(t)$ is the weight at each time step $t$, and $L_{\alpha_{k}}$ denotes the pinball loss function. Empirically, we set $K=9$ and use quantile levels of $\alpha=0.1,0.2,\ldots,0.9.$
D.2 Evaluation Details
D.2.1 Metric
@@
$\text{MSE}=\frac{1}{H}\sum_{i=1}^{H}(x_{i}-\hat{x})^{2},\quad\text{MAE}=\frac{1}{H}\sum_{i=1}^{H}\lvert x_{i}-\hat{x}\rvert$ (10)
where $x_{i}$ is the ground truth and $\hat{x}_{i}$ s the prediction for the $i$-th future time point.
D.2.2 Stride
Due to the significant inference latency of our baseline, Chronos (Ansari et al., 2024), we set the evaluation stride to 96 to enhance evaluation efficiency without compromising fairness. This setting is consistent with the protocols of Moirai (Woo et al., 2024), Moirai-MoE (Liu et al., 2024b), and the GIFT-Eval benchmark (Aksu et al., 2024), all of which set the stride equal to the prediction length. Additionally, the original Chronos paper evaluates the model on only a single window per dataset. Therefore, we wish to emphasize a critical point: while the specific numerical results would likely vary with a different stride, our chosen protocol is applied uniformly to all models under evaluation, ensuring a fair and equitable comparison.
D.3 Mixture-of-Size Encoder
D.3.1 Auxiliary-Loss-Free Load Balancing
In this section, we elaborate on the details of bias previously introduced in Section 3.2. In order to ensure that different experts are adequately trained and to control the distribution ratios of various patch sizes, we employ an Auxiliary-Loss-Free Load Balancing method similar to that used in DeepSeek-V3 (Liu et al., 2024a). Specifically, we sum the normalized gating value of the $i$-th expert activated throughout the entire sequence to obtain $L_{i}$:
$L_{i}=\sum_{n=1}^{N}\alpha_{n,i},$ (11)
where
$\alpha_{n,i}=\frac{g_{n,i}}{\sum_{j=1}^{S}g_{n,j}}.$ (12)
We define a target load distribution $\tau=(\tau_{1},\tau_{2},\ldots,\tau_{S})$, where $\tau_{i}$ is the desired proportion of the total load for expert $i$, satisfying $\sum_{j=1}^{S}\tau_{j}=1$. Next, we update the bias term $\boldsymbol{b}_{i}$ using $L_{i}$ and $\tau$:
$\boldsymbol{b}_{i}\leftarrow\boldsymbol{b}_{i}+\eta_{b}\cdot\frac{\tau_{i}\cdot\sum_{j=1}^{S}L_{j}-L_{i}}{\sum_{j=1}^{S}L_{j}},$ (13)
where $\eta_{b}$ is a hyper-parameter to govern the magnitude of this adjustment and is referred to as the bias update speed.
In practical implementation, $\boldsymbol{b}_{i}$ is updated once per batch, i.e., $L_{i}$ denotes the sum of the normalized gating values for the $i$-th expert across all sequences within the entire batch. At the end of each step, the bias term $\boldsymbol{b}_{i}$ for each expert is adjusted. This dynamic adjustment of $\boldsymbol{b}_{i}$ aims to balance the workload across the experts according to the desired distribution by influencing future top-K selections, while also steering the patches-to-expert affinity scores $s^{\prime}_{n,i}$ in subsequent batches, ensuring that the actual load distribution $L_{i}$ progressively aligns with the target distribution $\tau_{i}$, thereby promoting balanced expert utilization over time.
D.3.2 Definition of the Expansion Function
@@
  • Top-$K$ selection: $K=2$.
Suppose for a specific segment $\mathbf{x}$, the router activates experts corresponding to granularity levels 1 and 3 (i.e., $\mathcal{I}_{\mathrm{valid}}=\{1,3\}$). The finest activated resolution is $P_{\min}=\min(32,128)=32$. The target aligned sequence length is $M=128/32=4$.
@@
D.4 Dynamic Rotary Position Embedding (DRoPE)
In this section, we provide a more detailed description of the DRoPE implementation discussed in Section 3.3. We first introduce the principles of Rotary Position Embedding (RoPE) (Su et al., 2024) in detail, and then provide corresponding supplementary information on the DRoPE implementation.
D.4.1 Details of RoPE
@@
Theorem D.1 (Periodic dependence of RoPE attention).
@@ D.4.2 Details of DRoPE
$\log\theta^{\prime(l)}_{\text{inst},d}=\gamma^{(l)}_{d}\odot\log\theta_{d}+\beta^{(l)}_{d}$, (30)
@@
0: Time series instance $X\in\mathbb{R}^{B\times L}$, mask $M\in\{0,1\}^{B\times L}$, batch size $B$, sequence length $L$, FFT feature dimension $w$, MLP network $f_{\text{MLP}}$
@@
D.4.3 Interpretation of DRoPE
From the theorem D.1, the base RoPE parameters $\{\theta_{d}\}_{d=0}^{D_{h}/2-1}$ define a fixed bank of sinusoidal kernels over relative positions. Each $\theta_{d}$ controls the frequency of one cosine kernel. This fixed frequencies works reasonably well when sequences share a homogeneous notion of temporal scale, but it can become suboptimal for a TSFM that must handle highly heterogeneous sampling intervals and spectral patterns across domains.
@@ D.4.3 Interpretation of DRoPE
$\log\theta^{\prime(l)}_{\text{inst},d}=\gamma^{(l)}_{d}\odot\log\theta_{d}+\beta^{(l)}_{d},\quad\theta^{\prime(l)}_{\text{inst},d}=\exp\big(\log\theta^{\prime(l)}_{\text{inst},d}\big)$, (32)
@@
Appendix E Additional Details of Experiment Setting
E.1 Pre-training Datasets
@@
We trained Kairos on the Predictability-Stratified Time Series (PreSTS) corpus, which consists of over 300B real-world time series observations from Chronos (Ansari et al., 2024) and Moirai (Woo et al., 2024) in conjunction with 15B synthetic time points. Following (Das et al., 2024), the training loader samples 80% real data and 20% synthetic data.
@@
Real-world data. The real-world datasets were stratified into five tiers based on their predictability. This hierarchical structure dictates the sampling probability during model training, assigning a higher likelihood of selection to datasets with greater predictability. Such a strategy ensures that the model is preferentially trained on high-quality data while preserving its capacity to predict corner cases. Specifically, Tier 1 comprises datasets characterized by pronounced periodicity and trends with low noise. Tier 2 contains datasets with similarly distinct patterns but high noise, whereas Tier 3 includes those with subtle trends and considerable noise. The remaining datasets were classified into Tiers 4 and 5, based on a composite assessment of their size and pattern regularity. The specific details of these datasets, categorized by their respective sampling frequencies, are presented in Tables 6-9.
@@
Synthetic data. We build a synthetic data generator that produces two distinct types of time series, each with a length of 4096. The first type consists of composite series, created by the additive combination of seasonal, trend, and noise components. For seasonality, we sample one or two components, where the primary period is drawn from {24, 48, 288, 360}, and a potential second period is a fixed seven-fold multiple of the first; the seasonal patterns manifest as either spike trains or smooth non-sinusoidal templates generated via interpolation. An optional trend is chosen from linear, exponential, or an ARIMA-like process derived from cumulating a stationary ARMA model’s output. High-probability white Gaussian noise is also added, with each series guaranteed to contain at least one seasonal or trend component and all magnitudes bounded for stability. The generation process for these composite series is formally detailed in Algorithm 2. Complementing these are idealized industrial signals, which simulate perfectly regular machine cycles. These feature a constant baseline with repeating events like trapezoidal spikes or inverted-U shaped dips, where the period, amplitude, and width of the events remain fixed across the entire series with no random jitter, as outlined in Algorithm 3. Synthetic dataset cases are provided in Appendix H.1.
@@
Distribution Analysis. To investigate whether the synthetic dataset supplements distributions absent in the real-world data, we adopted the statistical methodology of Toto (Cohen et al., 2025). We employed the ARCH-LM Statistic and Spectral Entropy to quantify time-varying volatility and information density (complexity), respectively. These metrics align closely with the design motivation for the Mixture-of-Size Encoding and DRoPE components of Kairos. As illustrated in Figures 6 (a) and (b), we observed that a significant portion of the real-world dataset exhibits ARCH-LM Statistic values approaching zero and Spectral Entropy values nearing one, indicating high time-varying volatility and complexity. In contrast, the synthetic dataset exhibits the opposite characteristics, exposing the model during training to time series with more constant volatility and greater regularity. This exposure enhances the model’s generalization capabilities. Furthermore, we statistically compared the periodicity of the real-world and synthetic datasets by plotting their respective Cumulative Distribution Functions (CDFs). As shown in Figure 6 (c), the dominant periods in the real-world dataset are heavily concentrated. This can cause the model to overfit to these specific periodicities and fail to learn generalizable periodic patterns. Conversely, the synthetic dataset presents a significantly smoother distribution of periods, enabling the model to generalize effectively across arbitrary periodicities. Consequently, the synthetic dataset effectively mitigates these biases present in the real-world data, thereby exposing the model to a more comprehensive and diverse set of distributions.
@@
E.2 Evaluation Datasets

### v3

@@ Appendix D Analysis of Multi-Patch Prediction
Unlike task-specific, non-autoregressive models such as TFT [22] and Informer [46], which predict all future time points simultaneously, their architecture confines them to forecasting a fixed length identical to the predefined prediction horizon used during training, precluding variable-length predictions during inference. Conversely, TSFMs typically employ an iterative process: they predict a segment of the sequence, append it to the historical data, and then forecast the subsequent segment, thereby enabling autoregressive predictions of arbitrary length. To mitigate the cumulative errors inherent in this autoregressive methodology, Kairos introduces forecast tokens designed to predict multiple patches within each autoregressive step.
As briefly mentioned in Section 3.4, the multi-token design of our Multi-Patch Decoder is strategically motivated by two limitations in predominant forecasting paradigms: (i) Error mitigation and efficiency: Compared to conventional single-step autoregressive models [4, 27], predicting $J$ patches in parallel improves computational efficiency and mitigates cumulative error by reducing iterative decoding steps. (ii) Horizon flexibility: Unlike models that generate a single larger patch [11], our approach offers superior granularity for variable prediction horizons. By assembling the forecast from the specific subset of tokens required, Kairos ensures high-fidelity alignment with the target length and avoids the imprecise truncation associated with oversized outputs.
To empirically validate this design and determine the optimal setting, we present a detailed analysis of the multi-patch prediction strategy. During the training phase, we experimented with a range of forecast patch numbers, specifically $J=\{1,2,4,6,12\}$. The corresponding evaluation results on the GIFT-Eval benchmark are illustrated in Figure 5. Our observations reveal that when $J=1$, which corresponds to the conventional approach of predicting a single patch [4, 27, 11], the model achieves optimal performance in short-term forecasting. However, this method necessitates multiple iterations of autoregressive prediction, leading to a significant degradation in performance for medium- and long-term forecasting.
@@
Appendix E Validation of Learned Segment-Level Adaptation
To explicitly demonstrate that the performance gains of the Mixture-of-Size Encoder stem from genuine segment-level adaptation rather than merely a generic local multi-scale inductive bias, we conduct two dedicated analyses: a structural comparison with a sequence-level multi-scale baseline, and a direct causal intervention on the learned routing decisions.
Advantage of Intra-Sequence Adaptation. We separate the benefits of local intra-sequence adaptation from global multi-scale availability. To achieve this, we replaced our Mixture-of-Size Encoder with the sequence-level multi-scale routing mechanism proposed in Pathformer [8]. While Pathformer adaptively selects patch sizes for each sequence, it applies these selected granularities uniformly across the entire input, ignoring intra-sequence heterogeneity. As shown in Table 5, this replacement degrades GIFT-Eval MASE from 0.738 to 0.759, underscoring the advantage of local intra-sequence adaptation over global sequence-level multi-scale modeling.
Causal Intervention on Routing Decisions. To more directly test whether the router makes meaningful local decisions, we performed causal interventions on the routing mechanism at inference time. We kept all model weights and architectural components unchanged, but actively manipulated the router’s outputs:
  • Uniform Weights: Assigning uniform weights across all granularities within each segment, which removes adaptive selection while preserving the multi-scale structure.
  • Shuffled Routing: Randomly shuffling the router-produced granularity weights within each segment, which disrupts the learned routing decisions.
Assigning uniform weights across all granularities degrades performance from 0.738 to 0.831, indicating that the gain cannot be explained by multi-scale modeling alone. When we further disrupt the learned routing policy by randomly shuffling the granularity weights, performance deteriorates much more sharply to 1.205. Taken together, these results provide direct intervention evidence that Kairos benefits not simply from a generic multi-scale gain, but from assigning granularities appropriately to each segment.
Table 5: Empirical validation of segment-level adaptation on GIFT-Eval. Results report the normalized MASE across prediction horizons and as an aggregate across all tasks.
Setting Short Medium Long AVG
[…]
Structural Replacement
   w/ Sequence-Level Routing (Pathformer) 0.730 0.782 0.819 0.759
Causal Interventions at Inference
   w/ Uniform Granularity Weights 0.856 0.780 0.819 0.831
   w/ Shuffled Routing Decisions 1.154 1.282 1.271 1.205
Appendix F Analysis of Inference Speed
@@
To provide a comprehensive view of our model’s efficiency, we benchmarked the single-batch inference speed of Kairos against several state-of-the-art models. The experimental setup involved an input sequence length of 2048 and a prediction horizon of 96 on a single NVIDIA TITAN RTX GPU. For TTM-Advanced [13], its maximum supported input length of 1536 was used.
@@
As shown in Table 6, Kairos-Base occupies a superior position on the efficiency-performance frontier. Its inference time (0.061s) is on the same order of magnitude as other highly efficient models like ChronosBolt and Moirai, yet it delivers significantly better zero-shot accuracy (0.738 MASE). Compared to high-capacity models like Toto and Chronos, Kairos is several orders of magnitude faster while maintaining higher forecasting precision.
@@
Table 6: Comparison of single-batch inference speeds and zero-shot performance on GIFT-Eval. All models were tested with an input length of 2048 and an output length of 96, except for TTM-Advanced (*), which used its maximum input length of 1536. Inference times are averaged per batch on a single GPU.
Model Inference Time (s) $\downarrow$ MASE $\downarrow$
TTM-Advanced* 0.009 0.812
Timer-XL 0.014 0.795
ChronosBolt-Base 0.055 0.781
[…]
Moirai-Large 0.070 0.765
Sundial-Base 0.099 0.750
Time-MoE-Large 0.148 0.755
TimesFM-2.0 0.202 0.760
Chronos-Large 4.901 0.876
Toto-Base 13.717 0.750
Appendix G Decoupling Architecture and Data Contributions
In this section, we provide a dedicated experimental analysis to explicitly decouple the performance contributions of our proposed PreSTS corpus from the architectural innovations of Kairos.
Ablation on Data Curation. We isolate the impact of our data curation components, specifically the predictability-based tier-stratified sampling and the inclusion of synthetic data. As shown in Table 7, removing the data curation components degrades the performance on the GIFT-Eval benchmark, confirming their utility in providing high-quality supervision. However, this degradation is notably smaller than the performance drop observed when removing the main architectural components. This indicates that while the PreSTS data curation strategy provides a complementary gain, the architectural design serves as the primary contributor to the overall performance of Kairos.
Table 7: Ablation study decoupling data curation and architectural components. Models were evaluated using the normalized MASE on GIFT-Eval.
Variant MASE $\downarrow$
Data Curation Ablations
   w/o tier-stratified sampling 0.748
   w/o tier-stratified sampling & synthetic data 0.755
Architectural Ablations
   w/o Multi-Patch Decoder 0.753
   w/o Mixture-of-Size Encoder 0.761
   w/o DRoPE 0.767
[…]
Matched-Data Comparison. To further disentangle the architecture from the pre-training data, we perform a matched-data comparison. Specifically, we train Kairos on the original Chronos corpus [4] and train the advanced baseline ChronosBolt on our proposed PreSTS corpus. The results are summarized in Table 8.
Table 8: Matched-data comparison on the GIFT-Eval benchmark separating architecture and training data contributions.
Model Params Training Data MASE $\downarrow$
[…]
Kairos 53M Chronos corpus 0.761
ChronosBolt 205M PreSTS 0.781
ChronosBolt 205M Chronos corpus 0.808
Chronos 709M Chronos corpus 0.870
These matched-data results demonstrate that the architectural contribution significantly outweighs the data-curation contribution in our setting:
[…]
  • Under the same Chronos corpus, the gap between Kairos and ChronosBolt is 0.047 (0.761 vs. 0.808), whereas for ChronosBolt the gain from replacing the Chronos corpus with PreSTS is 0.027 (0.808 to 0.781). This suggests that, in our experiments, architectural innovation is the primary source of improvement, while PreSTS provides an additional but smaller gain.
Appendix H Implementation Details
H.1 Training Details
H.1.1 Model Configurations
@@
H.1.2 Loss Function
To better accommodate varied forecast horizons, and following the methodology of ElasTST [44], we build upon the standard quantile loss by assigning distinct weights to each timestep, such that earlier predictions are given greater importance. The model’s parameters are optimized by minimizing the quantile loss with weight decay, formulated as:
$\mathcal{L}=\frac{1}{B}\sum_{b=1}^{B}\sum_{t=1}^{H}\frac{1}{Q}\sum_{k=1}^{Q}\omega(t)L_{\alpha_{k}}(y_{b,t},q_{b,t}(\alpha_{k})),$ (8)
$\omega(t)=\frac{1}{H}(\ln(H)-\ln(t^{\prime})),\quad L_{\alpha}(y,q)=(\alpha-\mathbf{1}_{\{y<q\}})(y-q),$ (9)
where $B,H,Q$ are the batch size, forecast horizon, and total number of quantiles, respectively; $y_{b,t}$ is the ground truth value, $q_{b,t}(\alpha_{k})$ is the $k$-th quantile forecast, $\omega(t)$ is the weight at each time step $t$, and $L_{\alpha_{k}}$ denotes the pinball loss function. To prevent the loss weight from evaluating to zero at the final prediction step ($t=H$), we evaluate the logarithmic decay using $t^{\prime}$, which maps the discrete timesteps $t\in\{1,\dots,H\}$ to an evenly spaced continuous sequence from $1+\epsilon_{1}$ to $H-\epsilon_{2}$ (specifically, $\epsilon_{1}=10^{-5}$ and $\epsilon_{2}=10^{-3}$ in our implementation). Empirically, we set $Q=9$ and use quantile levels of $\alpha\in\{0.1,0.2,\dots,0.9\}$.
H.2 Evaluation Details
H.2.1 Metric
@@
$\text{MSE}=\frac{1}{H}\sum_{i=1}^{H}(x_{i}-\hat{x_{i}})^{2},\quad\text{MAE}=\frac{1}{H}\sum_{i=1}^{H}\lvert x_{i}-\hat{x_{i}}\rvert$ (10)
where $x_{i}$ is the ground truth and $\hat{x}_{i}$ is the prediction for the $i$-th future time point.
H.2.2 Stride
Due to the significant inference latency of our baseline, Chronos [4], we set the evaluation stride to 96 to enhance evaluation efficiency without compromising fairness. This setting is consistent with the protocols of Moirai [40], Moirai-MoE [25], and the GIFT-Eval benchmark [3], all of which set the stride equal to the prediction length. Additionally, the original Chronos paper evaluates the model on only a single window per dataset. Therefore, we wish to emphasize a critical point: while the specific numerical results would likely vary with a different stride, our chosen protocol is applied uniformly to all models under evaluation, ensuring a fair and equitable comparison.
H.3 Mixture-of-Size Encoder
H.3.1 Auxiliary-Loss-Free Load Balancing
In this section, we elaborate on the details of bias previously introduced in Section 3.2. In order to ensure that different experts are adequately trained and to control the distribution ratios of various patch sizes, we employ an Auxiliary-Loss-Free Load Balancing method similar to that used in DeepSeek-V3 [24].
Specifically, we compute the empirical load $L_{i}$ for the $i$-th expert by summing its normalized routing weights across all $N$ segments of all $B$ sequences within a training batch:
$L_{i}=\sum_{b=1}^{B}\sum_{n=1}^{N}\tilde{s}_{b,n,i},$ (11)
where $\tilde{s}_{b,n,i}$ denotes the routing weights for the $i$-th expert on the $n$-th segment of the $b$-th sequence, which directly corresponds to the definition of $\tilde{s}_{i}$ in Section 3.2.
We define a target load distribution $\tau=(\tau_{1},\dots,\tau_{G+Z})$, where $\tau_{i}$ is the desired proportion of the total load for expert $i$, satisfying $\sum_{j=1}^{G+Z}\tau_{j}=1$. At the end of each training step, we update the bias term $\boldsymbol{b}_{i}$ using the empirical load $L_{i}$ and the target $\tau$:
$\boldsymbol{b}_{i}\leftarrow\boldsymbol{b}_{i}+\eta_{b}\cdot\frac{\tau_{i}\cdot\sum_{j=1}^{G+Z}L_{j}-L_{i}}{\sum_{j=1}^{G+Z}L_{j}},$ (12)
where $\eta_{b}$ is a hyper-parameter governing the magnitude of this adjustment, referred to as the bias update speed.
This dynamic adjustment of $\boldsymbol{b}_{i}$ aims to balance the workload across the experts according to the desired distribution by influencing future top-K selections, while also steering the patches-to-expert affinity scores $s^{\prime}_{n,i}$ in subsequent batches, ensuring that the actual load distribution $L_{i}$ progressively aligns with the target distribution $\tau_{i}$, thereby promoting balanced expert utilization over time.
H.3.2 Definition of the Expansion Function
@@
  • Top-$K$ selection: $K=3$.
Suppose for a specific segment $\mathbf{x}$, the router activates experts corresponding to granularity levels 1 and 3, alongside one null expert (i.e., the selected set is $\{1,3,\text{null}\}$ while $\mathcal{I}_{\mathrm{valid}}=\{1,3\}$). The finest activated resolution is $P_{\min}=\min(32,128)=32$. The target aligned sequence length is $M=128/32=4$.
@@
H.4 Dynamic Rotary Position Embedding (DRoPE)
In this section, we provide a more detailed description of the DRoPE implementation discussed in Section 3.3. We first introduce the principles of Rotary Position Embedding (RoPE) [35] in detail, and then provide corresponding supplementary information on the DRoPE implementation.
H.4.1 Details of RoPE
@@
Theorem H.1 (Periodic dependence of RoPE attention).
@@ H.4.2 Details of DRoPE
$\log\theta^{\prime(l)}_{\text{inst},d}=\gamma^{(l)}_{d}\cdot\log\theta_{d}+\beta^{(l)}_{d}$, (29)
@@
0: Time series instance $X\in\mathbb{R}^{B\times L}$, mask $M\in\{0,1\}^{B\times L}$, batch size $B$, sequence length $L$, FFT feature dimension $\omega$, MLP network $f_{\text{MLP}}$
@@
H.4.3 Interpretation of DRoPE
From the theorem H.1, the base RoPE parameters $\{\theta_{d}\}_{d=0}^{D_{h}/2-1}$ define a fixed bank of sinusoidal kernels over relative positions. Each $\theta_{d}$ controls the frequency of one cosine kernel. This fixed frequencies works reasonably well when sequences share a homogeneous notion of temporal scale, but it can become suboptimal for a TSFM that must handle highly heterogeneous sampling intervals and spectral patterns across domains.
@@ H.4.3 Interpretation of DRoPE
$\log\theta^{\prime(l)}_{\text{inst},d}=\gamma^{(l)}_{d}\cdot\log\theta_{d}+\beta^{(l)}_{d},\quad\theta^{\prime(l)}_{\text{inst},d}=\exp\big(\log\theta^{\prime(l)}_{\text{inst},d}\big)$, (31)
@@
Appendix I Additional Details of Experiment Setting
I.1 Pre-training Datasets
@@
We trained Kairos on the Predictability-Stratified Time Series (PreSTS) corpus, which consists of over 300B real-world time series observations from Chronos [4] and Moirai [40] in conjunction with 15B synthetic time points. Following [11], the training loader samples 80% real data and 20% synthetic data.
@@
Real-world data. The real-world datasets were stratified into five tiers based on their predictability. This hierarchical structure dictates the sampling probability during model training, assigning a higher likelihood of selection to datasets with greater predictability. Such a strategy ensures that the model is preferentially trained on high-quality data while preserving its capacity to predict corner cases. Specifically, Tier 1 comprises datasets characterized by pronounced periodicity and trends with low noise. Tier 2 contains datasets with similarly distinct patterns but high noise, whereas Tier 3 includes those with subtle trends and considerable noise. The remaining datasets were classified into Tiers 4 and 5, based on a composite assessment of their size and pattern regularity. The specific details of these datasets, categorized by their respective sampling frequencies, are presented in Tables 10-13.
@@
Synthetic data. We build a synthetic data generator that produces two distinct types of time series, each with a length of 4096. The first type consists of composite series, created by the additive combination of seasonal, trend, and noise components. For seasonality, we sample one or two components, where the primary period is drawn from {24, 48, 288, 360}, and a potential second period is a fixed seven-fold multiple of the first; the seasonal patterns manifest as either spike trains or smooth non-sinusoidal templates generated via interpolation. An optional trend is chosen from linear, exponential, or an ARIMA-like process derived from cumulating a stationary ARMA model’s output. High-probability white Gaussian noise is also added, with each series guaranteed to contain at least one seasonal or trend component and all magnitudes bounded for stability. The generation process for these composite series is formally detailed in Algorithm 2. Complementing these are idealized industrial signals, which simulate perfectly regular machine cycles. These feature a constant baseline with repeating events like trapezoidal spikes or inverted-U shaped dips, where the period, amplitude, and width of the events remain fixed across the entire series with no random jitter, as outlined in Algorithm 3. Synthetic dataset cases are provided in Appendix L.1.
@@
Distribution Analysis. To investigate whether the synthetic dataset supplements distributions absent in the real-world data, we adopted the statistical methodology of Toto [10]. We employed the ARCH-LM Statistic and Spectral Entropy to quantify time-varying volatility and information density (complexity), respectively. These metrics align closely with the design motivation for the Mixture-of-Size Encoding and DRoPE components of Kairos. As illustrated in Figures 6 (a) and (b), we observed that a significant portion of the real-world dataset exhibits ARCH-LM Statistic values approaching zero and Spectral Entropy values nearing one, indicating high time-varying volatility and complexity. In contrast, the synthetic dataset exhibits the opposite characteristics, exposing the model during training to time series with more constant volatility and greater regularity. This exposure enhances the model’s generalization capabilities. Furthermore, we statistically compared the periodicity of the real-world and synthetic datasets by plotting their respective Cumulative Distribution Functions (CDFs). As shown in Figure 6 (c), the dominant periods in the real-world dataset are heavily concentrated. This can cause the model to overfit to these specific periodicities and fail to learn generalizable periodic patterns. Conversely, the synthetic dataset presents a significantly smoother distribution of periods, enabling the model to generalize effectively across arbitrary periodicities. Consequently, the synthetic dataset effectively mitigates these biases present in the real-world data, thereby exposing the model to a more comprehensive and diverse set of distributions.
@@
I.2 Compared Methods
We compare Kairos with several state-of-the-art models, including TSFMs such as Sundial [28], Toto [10], YingLong [38], VisionTS [7], Time-MoE [34], ChronosBolt [4], TTM [13], Moirai [40], Timer-XL [27], TimesFM-2.0 [11], Chronos [4], and advanced full-shot deep learning models, including PatchTST [29], DLinear [42], iTransformer [26], Pathformer [8], and TimesNet [41].
I.3 Evaluation Datasets